1. WACC5% YearC0 Cash $1000 $250 $200 $200 $500 $250 $200 $250 $200 Flows 4 5 5 3 3 5 Find A. PV B. FV C. MIRR D. IRR

Please show your work
Answer #1

Please refer to the table below. In the second column I have shown how each row has been calculated:

Part (A) PV and part (B) FV have been calculated in the workings below:

WACC

R

5%

Year, N

0

1

2

3

4

5

6

7

8

Cash flows

C

(1,000.00)

250.00

200.00

200.00

500.00

250.00

200.00

250.00

200.00

PV Factor

(1+R)-N

1.0000

0.9524

0.9070

0.8638

0.8227

0.7835

0.7462

0.7107

0.6768

PV of Cash flows

C x PV Factor

(1,000.00)

238.10

181.41

172.77

411.35

195.88

149.24

177.67

135.37

PV

Sum of all PV of cash flows

661.78

FV Factor

(1+R)N

1.4775

1.4071

1.3401

1.2763

1.2155

1.1576

1.1025

1.0500

1.0000

FV of Cash flows

C x FV factor

(1,477.46)

351.78

268.02

255.26

607.75

289.41

220.50

262.50

200.00

FV

Sum of all FV of cash flows

977.75

Part (C) MIRR

FV of all future positive cash flows = FV of all cash flows - FV of initial investment = 977.75 (as calculated in Part (B)) - Initial investment x FV factor = 977.75 - [-1,000 x FV factor at year 0] = 977.75 - (-1,000) x 1.4775 = 2,455.21

MIRR

N = period = 8 years

FVofall futur epositivecashf lows Initialinvestment

= = (15021)1/8-1 1,000 = 11.88%

Part (D)

IRR

We will have no other option but to use the excel function called IRR to calculate the IRR

IRR = IRR(Cash flow 0, Cash flow 1,....Cash flow 8) = IRR(-1000, 250,200,....,200) = 19.59%